OpenStudy (anonymous):

Tell whether the sequence 1/3, 0, 1, -2… is arithmetic, geometric, or neither. Find the next three terms of the sequence. A. neither; 7, -20, 61 B. geometric;7, -20, 61 C. arithmetic; -13, 113, 3 D. geometric;-313, -559, -9727

OpenStudy (amorfide):

arithmetic sequences are sequences in which you add a positive or negative number to get from one number to the next geometric is where you multiply by a common number known as the ratio to get from one number to the next

OpenStudy (anonymous):

I know, but I'm not sure how they started with a positive fraction for this one, and ended with a negative whole number

OpenStudy (anonymous):

\[(-10+9 n-2 n^2)/(-13+4 n)\]

OpenStudy (anonymous):

@optiquest I'm not sure what that equation has to do with this?

OpenStudy (anonymous):

what equation are you using to find each term

OpenStudy (amorfide):

geometric series where a is the first number in sequence, r is the ratio a, ar, ar^2 to get the ratio you do the current term divided by the term before it ar/a = r ar^2/ar=r this tells you the common ratio, if it is the same throughout, then it is geometric arithmetic where a is the first number, d is the number you add on a, a+d, a+d+d, so we have a, a+d, a+2d to find out d we do the current term subtract the term before it a+d - a=d a+2d - (a+d)=d

OpenStudy (amorfide):

if you cant find a common difference, or common ratio, then it is neither

OpenStudy (anonymous):

Well. it can't be geometric because if it was, anything after 0 would still be 0..

OpenStudy (anonymous):

And for being arithmetic, I don't think it could be that, because its not a consistent change..

OpenStudy (amorfide):

then you got your answer

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

This is for a different question, but if you're multiplying to get to the next term, that's geometric? @amorfide

OpenStudy (amorfide):

aslong as it is multiplying by the same number, to get from one number to the next yes

OpenStudy (amorfide):

a, ar, ar*r, ar*r*r

OpenStudy (anonymous):

Okay, thank you!

OpenStudy (amorfide):

you can't have a. ar. arb, arbc where r, b and c are different multipliers

OpenStudy (anonymous):

So if it's 3, 12, 36.. That would be geometric, or no..?

OpenStudy (amorfide):

no

OpenStudy (amorfide):

21/3=4 36/12=3

OpenStudy (anonymous):

Oh.. but it's being multiplied by 4 each time..

OpenStudy (amorfide):

no it is not

OpenStudy (amorfide):

3r=12 r=4 12r=36 r=3

OpenStudy (anonymous):

I thought you were supposed to start from the beginning to find a pattern? That's what I've always been taught..

OpenStudy (amorfide):

yes we did start from the beginning

OpenStudy (amorfide):

your sequence is 3,12,36

OpenStudy (amorfide):

geometric series is a, ar, ar^2

OpenStudy (amorfide):

a=3

OpenStudy (anonymous):

I see what I did wrong now.. But either way this sequence isn't going to be geometric right?

OpenStudy (amorfide):

3r=12 12r=36 you get different value for r

OpenStudy (amorfide):

it is only geometric if you get the same r value

OpenStudy (anonymous):

Okay, thank you